Horizontal square shaped platform of side 10 m on the ground has initially sides coinciding with x-axis and y-axis as shown. A platform starts moving along x-axis with a velocity 10 m/s. A person standing on the platform starts moving on platform towards diagonally opposite corner as shown, with a velocity 5​√2m/s. A stationary person standing at origin (Not on the platform) throws a ball in air so that the ball hits the person on platform when it is exactly at corner of the platform. Find the velocity of the ball?

1
Expert's answer
2021-08-31T08:59:02-0400

"v_1=10,~v=5\sqrt 2,"

"x=10,~d=10\sqrt 2,"

"t=\frac dv=2,"

"l=v_1t+x=30,"

"y=x=10,"

"r=\sqrt{l^2+y^2}=10\sqrt{10},"

"u=\frac rt=5\sqrt{10}~\frac ms."